a plane is flying 6 miles above the surface of the earth. A passenger looks out of the window to the distant horizon. On a clear day how far is the horizon to the nearest mile? assume the radius of the earth is 3960 miles
Answer by algebrahouse.com(1659)

A right triangle is formed with 6 miles above the earth and 3,960 mile radius
being the legs of the right triangle. The distance to the horizon is the
hypotenuse.

a² + b² = c² {the Pythagorean Theorem}
a and b are the legs, c is the hypotenuse

3960² + 6² = c² {substituted into the Pythagorean Theorem}
15,681,600 + 36 = c² {evaluated exponents}
15,681,636 = c² {added}
c ≈ 3,960.005 {took square root of each side and rounded}

The horizon is approximately 3,960 miles away

For more help from me, visit:
www.algebrahouse.com